# Monotonicity of the ultrafilter number function

### v3

@@ 4. Models with the failure of monotonicity 1
The next lemma might be a kind of folklore.
@@
Suppose not. Take an $\mathrm{Add}(\delta,\mu)$-name $\dot{\mathcal{U}}$ for a uniform ultrafilter over $\kappa$ with character $\nu<\mu$, and $\mathrm{Add}(\omega,\mu)$-names $\{\dot{A}_{\alpha}\mid\alpha<\nu\}$ of a base for $\mathcal{U}$. Fix a sufficiently large regular cardinal $\theta$, and take $M\prec H_{\theta}$ such that $\nu+\kappa\subseteq M$ and $\left|{M}\right|<\mu$. By the $\kappa^{\prime}$-c.c. of $\mathrm{Add}(\delta,\mu)$, we may assume that for every $\alpha<\nu$, $\dot{A}_{\alpha}$ is an $\mathrm{Add}(\delta,M\cap\mu)$-name. Take a $(V,\mathrm{Add}(\delta,\mu))$-generic $G$, and let $H=G\cap\mathrm{Add}(\delta,M\cap\mu)$, which is $(V,\mathrm{Add}(\delta,M\cap\mu))$-generic. We know $(\dot{A}_{\alpha})_{G}=(\dot{A}_{\alpha})_{H}$, hence $(\dot{A}_{\alpha})_{G}\in V[H]$. By the standard genericity argument, the forcing $\mathrm{Add}(\delta,\mu\setminus M)$ over $V[H]$ adds a new set $B\in[\kappa]^{\kappa}$ such that $(\dot{A}_{\alpha})_{G}\nsubseteq B$ and $(\dot{A}_{\alpha})_{G}\nsubseteq\kappa\setminus B$ for every $\alpha<\nu$, this contradicts to that $\{(\dot{A}_{\alpha})_{G}\mid\alpha<\nu\}$ is a base for $\mathcal{U}$. ∎
@@
Suppose not. Take an $\mathrm{Add}(\delta,\mu)$-name $\dot{\mathcal{U}}$ for a uniform ultrafilter over $\kappa$ with character $\mu$, and $\mathrm{Add}(\delta,\mu)$-names $\{\dot{A}_{\alpha}\mid\alpha<\mu\}$ of a base for $\mathcal{U}$.
@@
    $\kappa\subseteq M_{i}$.
@@
For $i<\kappa$, since $\mathrm{Add}(\delta,\mu)$ has the $\kappa^{\prime}$-c.c., we may assume that for every $\alpha\in M_{i}\cap\mu$, $\dot{A}_{\alpha}$ is an $\mathrm{Add}(\delta,M_{i}\cap\mu)$-name. Fix a partition $\{C_{i}\mid i<\kappa\}$ of $\kappa$ with $\left|{C_{i}}\right|=\delta$. Fix also a bijection $\pi_{i}:\delta\to C_{i}$. Take a sequence $\langle{\beta_{i}\mid i<\kappa}\rangle$ in $\mu$ such that $\beta_{i}\notin M_{i}\cap\mu$ for $i<\kappa$.
Now take a $(V,\mathrm{Add}(\delta,\mu))$-generic $G$, and for $\gamma<\mu$, let $r_{\gamma}$ be the $\gamma$-th generic subset of $\delta$ induced by $G$. Since $\beta_{i}\notin M_{i}$, we know that $r_{\beta_{i}}$ is $V[G\cap\mathrm{Add}(\delta,M_{i}\cap\mu)]$-generic, hence we have that $s_{i}=\pi_{i}``r_{\beta_{i}}$ has cardinality $\delta$. Let $B=\bigcup_{i<\kappa}s_{i}$. We know $B\in[\kappa]^{\kappa}$. We prove that $(\dot{A}_{\alpha})_{G}\nsubseteq B$ for $\alpha<\mu$. Fix $\alpha<\mu$, and take a condition $p\in G$. Pick a large $i<\kappa$ such that $\alpha,p\in M_{i}$. Let $H=G\cap\mathrm{Add}(\delta,M_{i}\cap\mu)$. We know $(\dot{A}_{\alpha})_{G}=(\dot{A}_{\alpha})_{H}$, hence $(\dot{A}_{\alpha})_{G}\in V[H]$. Since $\bigcup_{j\leq i}C_{j}$ has cardinality $<\kappa$ but $(\dot{A}_{\alpha})_{G}$ has $\kappa$, we can take $k>i$ and $\xi\in(\dot{A}_{\alpha})_{G}\cap C_{k}$. By the definition of $s_{k}$, there is a condition $q\in\mathrm{Add}(\delta,\mu\setminus M_{i})$ such that, in $V[H]$, $q$ forces that $\xi\notin\dot{s}_{k}$, where $\dot{s}_{k}$ is a canonical name for $s_{k}$. Thus, in $V$, there is some $p^{\prime}\leq p$ in $\mathrm{Add}(\delta,\mu)$ such that $p^{\prime}\Vdash\dot{A}_{\alpha}\nsubseteq\dot{B}$, where $\dot{B}$ is a canonical name for $B$. A similar argument shows that $(\dot{A}_{\alpha})_{G}\nsubseteq\kappa\setminus B$ for $\alpha<\mu$. These means that $B,\kappa\setminus B\notin(\dot{\mathcal{U}})_{G}$, this is a contradiction. ∎
@@
We does not need this lemma if $\mathord{\mathrm{cf}}(\mathfrak{u}(\kappa))\neq\mathord{\mathrm{cf}}(\kappa)$ is provable from $\mathsf{ZFC}$, and it is an open problem asked in [6].
@@
We note that $\omega_{1}$ can be replaced by any cardinal $<\kappa$ with uncountable cofinality.
@@
By using Fact 2.6 we have:
@@
If $\kappa$ is strongly compact, then every regular cardinal $\lambda\geq\kappa$ has a $\kappa$-complete uniform ultrafilter. By this fact, we have:
@@ 5. Easy $\mathsf{ZFC}$-results
The following lemmas will be used frequently.
@@
Since $\mathcal{U}$ is $\lambda$-decomposable, there is $f:\kappa\to\lambda$ such that $f_{*}(\mathcal{U})$ is a uniform ultrafilter over $\lambda$. If $\mathcal{G}$ is a base for $\mathcal{U}$, then $\{f``A\mid A\in\mathcal{G}\}$ generates $f_{*}(\mathcal{U})$; For $A\in\mathcal{G}$, since $A\subseteq f^{-1}(f``A)$, we have $f``A\in f_{*}(\mathcal{U})$. For the converse, if $X\in f_{*}(\mathcal{U})$, pick $A\in\mathcal{G}$ with $A\subseteq f^{-1}(X)$. Then $f``A\subseteq X$. These show that $\{f``A\mid A\in\mathcal{G}\}$ generates $f_{*}(\mathcal{U})$. Therefore $\mathfrak{u}(\lambda)\leq\chi(f_{*}(\mathcal{U}))\leq\chi(\mathcal{U})$. ∎
Lemma 5.2.
@@
Proof.
Take any uniform ultrafilter $\mathcal{U}$ over $\kappa$ with $\chi(\mathcal{U})=\mathfrak{u}(\kappa)$. If it is $\lambda$-decomposable, then $\mathfrak{u}(\lambda)\leq\chi(\mathcal{U})=\mathfrak{u}(\kappa)$ by Lemma 5.1, this contradicts to the assumption. ∎
@@
If not, then $\kappa$ has an $\omega$-indecomposable uniform ultrafilter $\mathcal{U}$ by Lemma 5.2. Then $\mathcal{U}$ is in fact $\sigma$-complete, so there is a measurable cardinal $\lambda\leq\kappa$. Hence we have $\mathfrak{u}(\omega)\leq 2^{\omega}<\lambda\leq\kappa\leq\mathfrak{u}(\kappa)$, this is a contradiction. ∎
@@
(2). By (1), we have $\mathfrak{u}(\lambda)\leq\mathfrak{u}(\lambda^{+})\leq\mathfrak{u}(\lambda^{++})\leq\cdots$, hence $\lambda^{+\omega}\leq\kappa$. ∎
@@ 6. Models with the failure of monotonicity 2
In Section 4, we constructed some models with $\lambda<\kappa$ but $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$. Our main tools were Raghavan and Shelah’s theorem 3.1 and Lemmas 4.1, 4.2. However, to use these tools, it is necessary that $\mathord{\mathrm{cf}}(\lambda)>\omega$, and we cannot apply our argument to the case $\mathord{\mathrm{cf}}(\lambda)=\omega$ easily. In this section, we show that it is possible that $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{\omega})$ for some $\kappa>\omega_{\omega}$.
@@
We claim that there is $f\in\mathcal{F}$ which is $<_{\mathcal{U}}$-minimal in $\mathcal{F}$, that is, if $g\in\mathcal{F}$ then $\{\alpha<\kappa\mid g(\alpha)\geq f(\alpha)\}\in\mathcal{U}$. Such a $<_{\mathcal{U}}$-minimal function is a require one; This can be verified as follows. It is easy to check that $f_{*}(\mathcal{U})$ is uniform. First note that for every $g\in\mathcal{F}$, since $g\circ f\in\mathcal{F}$ and $f\leq_{U}g\circ f$, if $g\leq\mathrm{id}$ then $\{\alpha<\kappa\mid g(\alpha)=\alpha\}\in f_{*}(\mathcal{U})$ Now take $C\subseteq\kappa$ such that $C\cap\kappa_{n}$ is a club in $\kappa_{n}$ for every $n<\omega$. Define $g:\kappa\to\kappa$ by $g(\alpha)=\sup(C\cap\alpha)$. Clearly $g\in\mathcal{F}$ and $g\leq\mathrm{id}$. Hence $X=\{\alpha<\kappa\mid g(\alpha)=\alpha\}\in f_{*}(\mathcal{U})$. By the definition of $g$, we have $X\subseteq(C\setminus(\min(C)+1))\cup\{\kappa_{n}\mid n<\omega\}$. On the other hand, since $f_{*}(\mathcal{U})$ is uniform, we have $\{\kappa_{n}\mid n<\omega\}\notin\mathcal{U}$, hence $C\in f_{*}(\mathcal{U})$.
We show that $\mathcal{F}$ has a $<_{\mathcal{U}}$-minimal element. Suppose to the contrary that such an element does not exist. By induction on $\xi<\lambda$, we take $f_{\xi}\in\mathcal{F}$ so that:
@@
The following can be seen as a diagonal version of Prikry and Silver’s theorem (Fact 2.4).
@@
    For every large regular cardinal $\theta$, the set $\{M\in[H_{\theta}]^{\lambda}\mid{}^{\omega}M\subseteq M$, and $\forall n<\omega(\sup(M\cap\kappa_{n})\in S_{n})\}$ is stationary in $[H_{\theta}]^{\lambda}$.
@@
For $n<\omega$ and $\alpha\in S_{n}$, fix a club $c_{\alpha}\subseteq\alpha$ with order type $\lambda$. Let $\langle{\xi_{i}^{\alpha}\mid i<\lambda}\rangle$ be the increasing enumeration of $c_{\alpha}$. The following is well-known, e.g., see Lemma 2.12 in Eisworth [5]:
Fact 6.3.
@@
Now suppose the contrary that there is a $\lambda$-indecomposable uniform ultrafilter $\mathcal{U}$ over $\kappa$. By Lemma 6.1, we may assume that for every $t\in\prod_{n<\omega}\kappa_{n}$, the set $\bigcup_{n<\omega}[t(n),\kappa_{n})$ is in $\mathcal{U}$. For $t\in\prod_{n<\omega}S_{n}$, define $\lambda(t)<\lambda$ as follows: By the choice of $\mathcal{U}$, we have that $X_{t}=\{\beta<\kappa\mid\beta>t(n(\beta))\}\in\mathcal{U}$. Define $G_{t}:X_{t}\to\lambda$ by $G_{t}(\beta)=F_{\beta}(t(n(\beta))$. Since $\mathcal{U}$ is $\lambda$-indecomposable, there is $\lambda(t)<\lambda$ such that $Y_{t}=\{\beta\in X_{t}\mid G_{t}(\beta)=F_{\beta}(t(n(\beta))<\lambda(t)\}\in\mathcal{U}$.
@@
For $t,t^{\prime}\in\prod_{n<\omega}S_{n}$, if $\{n<\omega\mid t(n)\neq t(n^{\prime})\}$ is cofinite, then there are infinitely many $n<\omega$ such that $\{\xi_{i}^{t(n)}\mid\lambda(t)<i<\lambda\}\cap\{\xi_{i}^{t^{\prime}(n)}\mid\lambda(t^{\prime})<i<\lambda\}=\emptyset$.
@@
We know $Y_{t}\cap Y_{t^{\prime}}\in\mathcal{U}$. Hence for a given $m<\omega$, we can find $\beta\in Y_{t}\cap Y_{t^{\prime}}$ such that $m<n(\beta)$, $t(n(\beta))\neq t^{\prime}(n(\beta))$, and $t(n(\beta)),t^{\prime}(n(\beta))<\beta$. By the choice of $F_{\beta}$, we have $\{\xi_{i}^{t(n(\beta))}\mid F_{\beta}(t(n(\beta)))<i<\lambda\}\cap\{\xi_{i}^{t^{\prime}(n(\beta))}\mid F_{\beta}(t^{\prime}(n(\beta)))<i<\lambda\}=\emptyset$. Because $\beta\in Y_{t}\cap Y_{t^{\prime}}$, we know $F_{\beta}(t(n(\beta)))<\lambda(t)$ and $F_{\beta}(t^{\prime}(n(\beta)))<\lambda(t^{\prime})$. Hence $\{\xi_{i}^{t(n(\beta))}\mid\lambda(t)<i<\lambda\}\cap\{\xi_{i}^{t^{\prime}(n(\beta))}\mid\lambda(t^{\prime})<i<\lambda\}=\emptyset$. ∎
@@
To show that the assumption of Proposition 6.2 is consistent, we use the forcing notion which adds a non-reflecting stationary subsets.
For a poset $\mathbb{P}$ and an ordinal $\alpha$, let $\Gamma_{\alpha}(\mathbb{P})$ denote the following two players game of length $\alpha$: At each inning, Players I and II choose conditions of $\mathbb{P}$ alternately with $p_{0}\geq q_{0}\geq p_{1}\geq q_{1}\geq\cdots$, but if $\beta<\alpha$ is limit, at the $\beta$-th inning, Player I does not move and only Player II choose a condition $q_{\beta}$ which is a lower bound of the partial play $\langle p_{\xi},q_{\zeta}\mid\xi,\zeta<\beta$, $\xi=0$ or $\rangle$ (if it is possible):
00 $1$cdots $\omega$ $\omega+1$cdots
I $p_{0}$ $p_{1}$cdots $p_{\omega+1}$cdots
II $q_{0}$ $q_{1}$cdots $q_{\omega}$ $q_{\omega+1}$cdots

### v4

@@ 4. Models with the failure of monotonicity 1
The next lemma is folklore.
@@
Suppose not. Take an $\mathrm{Add}(\delta,\mu)$-name $\dot{\mathcal{U}}$ for a uniform ultrafilter over $\kappa$ with character $\nu<\mu$, and $\mathrm{Add}(\delta,\mu)$-names $\{\dot{A}_{\alpha}\mid\alpha<\nu\}$ of a base for $\mathcal{U}$. Fix a sufficiently large regular cardinal $\theta$, and take $M\prec H_{\theta}$ such that $\nu+\kappa\subseteq M$, $\{\dot{A}_{\alpha}\mid\alpha<\nu\}\in M$, and $\left|{M}\right|<\mu$. By the $\kappa^{\prime}$-c.c. of $\mathrm{Add}(\delta,\mu)$, we may assume that for every $\alpha<\nu$, $\dot{A}_{\alpha}$ is an $\mathrm{Add}(\delta,M\cap\mu)$-name. Take a $(V,\mathrm{Add}(\delta,\mu))$-generic $G$, and let $H=G\cap\mathrm{Add}(\delta,M\cap\mu)$, which is $(V,\mathrm{Add}(\delta,M\cap\mu))$-generic. We know $(\dot{A}_{\alpha})_{G}=(\dot{A}_{\alpha})_{H}$, hence $(\dot{A}_{\alpha})_{G}\in V[H]$. $\mathrm{Add}(\delta,\mu\setminus M)$ is the quotient forcing from $V[H]$ to a generic extension of $\mathrm{Add}(\delta,\mu)$. Hence, by the standard genericity argument, forcing $\mathrm{Add}(\delta,\mu\setminus M)$ over $V[H]$ adds a new set $B\in[\kappa]^{\kappa}$ such that $(\dot{A}_{\alpha})_{G}\nsubseteq B$ and $(\dot{A}_{\alpha})_{G}\nsubseteq\kappa\setminus B$ for every $\alpha<\nu$, this contradicts to that $\{(\dot{A}_{\alpha})_{G}\mid\alpha<\nu\}$ is a base for $\mathcal{U}$. ∎
@@
Suppose not. Take an $\mathrm{Add}(\delta,\mu)$-name $\dot{\mathcal{U}}$ for a uniform ultrafilter over $\kappa$ with character $\mu$, and $\mathrm{Add}(\delta,\mu)$-names $\{\dot{A}_{\alpha}\mid\alpha<\mu\}$ of a base for $\mathcal{U}$. We may assume that each $\dot{A}_{\alpha}$ is a nice name, that is, there is a family $\{J_{k}\mid k<\kappa\}$ of antichains in $\mathrm{Add}(\delta,\mu)$ such that $\dot{A}_{\alpha}=\bigcup_{k<\kappa}(\{\check{k}\}\times J_{k})$
@@
    $\kappa+\kappa^{\prime}\subseteq M_{i}$.
@@
For $i<\kappa$, since $\mathrm{Add}(\delta,\mu)$ has the $\kappa^{\prime}$-c.c., we may assume that for every $\alpha\in M_{i}\cap\mu$, $\dot{A}_{\alpha}$ is an $\mathrm{Add}(\delta,M_{i}\cap\mu)$-name; $\dot{A}_{\alpha}$ is of the form $\bigcup_{k<\kappa}(\{\check{\alpha}\}\times J_{k})$ for some antichains $\{J_{k}\mid k<\kappa\}\in M_{\alpha}$. Since $\kappa,\kappa^{\prime}\subseteq M_{\alpha}$ and $\left|{J_{k}}\right|<\kappa^{\prime}$, we have $\dot{A}_{\alpha}\subseteq M_{\alpha}$, hence it can be seen as an $\mathrm{Add}(\delta,M_{i}\cap\mu)$-name.
Fix a partition $\{C_{i}\mid i<\kappa\}$ of $\kappa$ with $\left|{C_{i}}\right|=\delta$. Fix also a bijection $\pi_{i}:\delta\to C_{i}$. Take a sequence $\langle{\beta_{i}\mid i<\kappa}\rangle$ in $\mu$ such that $\beta_{i}\notin M_{i}\cap\mu$ for $i<\kappa$.
Now take a $(V,\mathrm{Add}(\delta,\mu))$-generic $G$, and for $\gamma<\mu$, let $r_{\gamma}$ be the $\gamma$-th generic subset of $\delta$ induced by $G$. Since $\beta_{i}\notin M_{i}$, we know that $r_{\beta_{i}}$ is $V[G\cap\mathrm{Add}(\delta,M_{i}\cap\mu)]$-generic, hence we have that $s_{i}=\pi_{i}``r_{\beta_{i}}$ has cardinality $\delta$. Let $B=\bigcup_{i<\kappa}s_{i}$. We know $B\in[\kappa]^{\kappa}$. We prove that $(\dot{A}_{\alpha})_{G}\nsubseteq B$ for $\alpha<\mu$. Fix $\alpha<\mu$, and take a condition $p\in G$. Pick a large $i<\kappa$ such that $\alpha,p\in M_{i}$. Let $H=G\cap\mathrm{Add}(\delta,M_{i}\cap\mu)$. We know $(\dot{A}_{\alpha})_{G}=(\dot{A}_{\alpha})_{H}$, hence $(\dot{A}_{\alpha})_{G}\in V[H]$. Since $\bigcup_{j\leq i}C_{j}$ has cardinality $\kappa$ but $(\dot{A}_{\alpha})_{G}$ has $\kappa$, we can take $k>i$ and $\xi\in(\dot{A}_{\alpha})_{G}\cap C_{k}$. By the definition of $s_{k}$, there is a condition $q\in\mathrm{Add}(\delta,\mu\setminus M_{i})$ such that, in $V[H]$, $q$ forces that $\check{\xi}\notin\dot{s}_{k}$, where $\dot{s}_{k}$ is a canonical name for $s_{k}$. Thus, in $V$, there is some $p^{\prime}\leq p$ in $\mathrm{Add}(\delta,\mu)$ such that $p^{\prime}\Vdash\dot{A}_{\alpha}\nsubseteq\dot{B}$, where $\dot{B}$ is a canonical name for $B$. A similar argument shows that $(\dot{A}_{\alpha})_{G}\nsubseteq\kappa\setminus B$ for $\alpha<\mu$. These means that $B,\kappa\setminus B\notin(\dot{\mathcal{U}})_{G}$, this is a contradiction. ∎
@@
We do not need this lemma if $\mathord{\mathrm{cf}}(\mathfrak{u}(\kappa))\neq\mathord{\mathrm{cf}}(\kappa)$ is provable from $\mathsf{ZFC}$, and it is an open problem asked by Garti and Shelah [8].
@@
We note that in fact $\mathfrak{u}(\kappa)=\mu$ holds in the resulting model of this proposition, and this was already remarked by Benhamou [3, Proposition 5.12].
@@
We note that $\omega_{1}$ can be replaced by any cardinal $\kappa$ with uncountable cofinality.
@@
By using Theorem 2.6 we have:
@@
If $\kappa$ is strongly compact, then every regular cardinal $\lambda\geq\kappa$ has a $\kappa$-complete uniform ultrafilter. By this theorem, we also have:
@@
$\omega_{1}$ in this theorem can be replaced by any regular uncountable cardinal $\mathrm{<}\kappa$. We also note that $\kappa^{+\omega_{1}}$ is not strong limit in the resulting model of this theorem. In Section 7, we prove that $\omega_{1}$ cannot be replaced by $\omega$, and if $\mathfrak{u}(\kappa^{+\omega_{1}+1})<\mathfrak{u}(\kappa^{+\omega_{1}})$, then $\kappa^{+\omega_{1}}$ is never strong limit.
@@ 5. Easy $\mathsf{ZFC}$-results
The following lemma is folklore but will be used frequently.
@@
Since $\mathcal{U}$ is $\lambda$-decomposable, by Lemma 2.8, there is $f:\kappa\to\lambda$ such that $f_{*}(\mathcal{U})$ is a uniform ultrafilter over $\lambda$ with $\chi(f_{*}(\mathcal{U}))\leq\chi(\mathcal{U})$. Thus we have $\mathfrak{u}(\lambda)\leq\chi(f_{*}(\mathcal{U}))\leq\chi(\mathcal{U})$. ∎
The next lemma is just the contraposition of Lemma 5.1.
Corollary 5.2.
@@
If not, then $\kappa$ has an $\omega$-indecomposable uniform ultrafilter $\mathcal{U}$ by Corollary 5.2. Then $\mathcal{U}$ is in fact $\sigma$-complete, so there is a measurable cardinal $\lambda\leq\kappa$. Hence we have $\mathfrak{u}(\omega)\leq 2^{\omega}<\lambda\leq\kappa\leq\mathfrak{u}(\kappa)$, this is a contradiction. ∎
@@
(2). By (1), we have $\mathfrak{u}(\lambda)\leq\mathfrak{u}(\lambda^{+})\leq\mathfrak{u}(\lambda^{++})\leq\cdots$, hence $\lambda^{+\omega}\leq\kappa$.
@@ 6. Models with the failure of monotonicity 2
In Section 4, we constructed some models with $\lambda<\kappa$ but $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$. Our main tools were Raghavan and Shelah’s theorem 3.1 and Lemma 4.1, 4.2. However, we cannot use these tools when $\mathord{\mathrm{cf}}(\lambda)=\omega$. In this section, we show that it is possible that $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{\omega})$ for some $\kappa>\omega_{\omega}$.
@@
We claim that there is $f\in\mathcal{F}$ which is $_{\mathcal{U}}-$minimal in $\mathcal{F}$, that is, if $g\in\mathcal{F}$ then $\{\alpha<\kappa\mid g(\alpha)\geq f(\alpha)\}\in\mathcal{U}$. Such a $_{\mathcal{U}}-$minimal function is as wanted; This can be verified as follows. Take $C\subseteq\kappa$ such that $C\cap\kappa_{n}$ is a club in $\kappa_{n}$ for every $n<\omega$. Define $g:\kappa\to\kappa$ by $g(\alpha)=\sup(C\cap\alpha)$. Clearly $g\in\mathcal{F}$ and $g\leq\mathrm{id}$. Since $g\circ f\in\mathcal{F}$ and $f\leq_{U}g\circ f$, we have $X=\{\alpha<\kappa\mid g(\alpha)=\alpha\}\in f_{*}(\mathcal{U})$. By the definition of $g$, we also have $X\subseteq(C\setminus(\min(C)+1))\cup\{\kappa_{n}\mid n<\omega\}$. On the other hand, since $f_{*}(\mathcal{U})$ is uniform, we have $\{\kappa_{n}\mid n<\omega\}\notin\mathcal{U}$, hence $C\in f_{*}(\mathcal{U})$.
We show that $\mathcal{F}$ has a $_{\mathcal{U}}-$minimal element. Suppose to the contrary that such an element does not exist. By induction on $\xi<\lambda$, we take $f_{\xi}\in\mathcal{F}$ so that:
@@
The following can be seen as a diagonal version of Prikry and Silver’s theorem (Theorem 2.4).
@@
    For every enough large regular cardinal $\theta$, the set $\{M\in[H_{\theta}]^{\lambda}\mid{}^{\omega}M\subseteq M$, and $\forall n<\omega(\sup(M\cap\kappa_{n})\in S_{n})\}$ is stationary in $H_{\theta}]^{\lambda}$.
@@
Note that the assumption (3) of this proposition is different from the notion of mutually stationary, which only requires the set $\{M\in[H_{\theta}]^{\lambda}\mid$ $\forall n<\omega(\sup(M\cap\kappa_{n})\in S_{n})\}$ is stationary in $H_{\theta}]^{\lambda}$; $M$ need not to be closed under $\omega$-sequences.
@@
For $n<\omega$ and $\alpha\in S_{n}$, fix a club $c_{\alpha}\subseteq\alpha$ with order type $\lambda$. Let $\langle{\xi_{i}^{\alpha}\mid i<\lambda}\rangle$ be the increasing enumeration of $c_{\alpha}$. For the next well-known theorem, we use the fact that the sets $S_{n}$’s are non-reflecting. See Lemma 2.12 in Eisworth [7] for the proof.
Theorem 6.3.
@@
Now suppose the contrary that there is a $\lambda$-indecomposable uniform ultrafilter $\mathcal{U}$ over $\kappa$, in particular it is not $(\omega,\lambda)$-regular. By Lemma 2.8 and 6.1, we may assume that for every $t\in\prod_{n<\omega}\kappa_{n}$, the set $\bigcup_{n<\omega}[t(n),\kappa_{n})$ is in $\mathcal{U}$. For $t\in\prod_{n<\omega}S_{n}$, define $\lambda(t)<\lambda$ as follows: By the choice of $\mathcal{U}$, we have that $X_{t}=\{\beta<\kappa\mid\beta>t(n(\beta))\}\in\mathcal{U}$. Define $G_{t}:X_{t}\to\lambda$ by $G_{t}(\beta)=F_{\beta}(t(n(\beta))$. Since $\mathcal{U}$ is $\lambda$-indecomposable, there is $\lambda(t)<\lambda$ such that $Y_{t}=\{\beta\in X_{t}\mid G_{t}(\beta)=F_{\beta}(t(n(\beta)))<\lambda(t)\}\in\mathcal{U}$.
@@
For $t,t^{\prime}\in\prod_{n<\omega}S_{n}$, if $\{n<\omega\mid t(n)\neq t^{\prime}(n)\}$ is cofinite, then there are infinitely many $n<\omega$ such that $\{\xi_{i}^{t(n)}\mid\lambda(t)<i<\lambda\}\cap\{\xi_{i}^{t^{\prime}(n)}\mid\lambda(t^{\prime})<i<\lambda\}=\emptyset$.
@@
We know that $Y_{t}\cap Y_{t^{\prime}}\in\mathcal{U}$. For a given $m<\omega$, since $\mathcal{U}$ is uniform and the set $\{n<\omega\mid t(n)\neq t(n^{\prime})\}$ is cofinite, we have that the set $\{\beta<\kappa\mid n(\beta)>m,t(n(\beta))\neq t^{\prime}(n(\beta))\}$ is in $\mathcal{U}$. Thus we can find $\beta\in Y_{t}\cap Y_{t^{\prime}}$ such that $m<n(\beta)$, $t(n(\beta))\neq t^{\prime}(n(\beta))$, and $t(n(\beta)),t^{\prime}(n(\beta))<\beta$. By the choice of $F_{\beta}$, we have $\{\xi_{i}^{t(n(\beta))}\mid F_{\beta}(t(n(\beta)))<i<\lambda\}\cap\{\xi_{i}^{t^{\prime}(n(\beta))}\mid F_{\beta}(t^{\prime}(n(\beta)))<i<\lambda\}=\emptyset$. Since $\beta\in Y_{t}\cap Y_{t^{\prime}}$, we know $F_{\beta}(t(n(\beta)))<\lambda(t)$ and $F_{\beta}(t^{\prime}(n(\beta)))<\lambda(t^{\prime})$. Hence $\{\xi_{i}^{t(n(\beta))}\mid\lambda(t)<i<\lambda\}\cap\{\xi_{i}^{t^{\prime}(n(\beta))}\mid\lambda(t^{\prime})<i<\lambda\}=\emptyset$. ∎
@@
To show that the assumption of Proposition 6.2 is consistent, we use a forcing notion which adds a non-reflecting stationary subset.
[…]
0 11 $\cdots$ $\omega$ $\omega+1$ $\cdots$
I $p_{0}$ $p_{1}$ $\cdots$ $p_{\omega+1}$ $\cdots$
II $q_{0}$ $q_{1}$ $\cdots$ $q_{\omega}$ $q_{\omega+1}$ $\cdots$